If 10.0 moles of O₂ are reacted with excess NO in the reaction below, and only 6.4 mol of NO₂ were collected, then what is the percent yield for the reaction?

2 NO (g) + O₂ (g) → 2 NO₂ (g)


1
Expert's answer
2022-02-16T23:16:02-0500

Moles of O2 = 10.0mol

Mole ratio of O2:NO2 = 1:2

Theoretical moles of NO2 "=\dfrac{2}{1}x10.0 mol" = 20.0mol

Experimental moles = 6.4mol


% yield "=\dfrac{experimental}{theoretical}x100"


"=\dfrac{6.4mol}{20.0mol}x100"


= 32%